MILLIMETER WAVE SPECTRUM OF CoCO(X^{2}Δ_{i}) IN THE GROUND AND VIBRATIONALLY EXCITED STATES

MASATO HAYASHI, \underline{TOSHIHIDE HIKIDA}, KENSUKE HARADA, KEIICHI TANAKA and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

Rotational spectrum of the CoCO radical generated by ultraviolet photolysis of Co(CO)_{3}NO was measured in the millimeter wave region to determine rotational and hyperfine interaction constants. Rotational transitions ranging from J = 29.5 − 28.5 to 34.5 − 33.5 were assigned in the Ω = 3/2 and Ω = 5/2 spin substates of the X^{2}Δ_{i} ground vibronic state as well as in the vibrationally excited states ν_{2}, ν_{3}, and 2ν_{2} of the Ω = 5/2 spin substate. Each rotational transition was split into 8 hyperfine components due to the ^{59}Co nucleus. Molecular constants, including the rotational constant B, centrifugal distortion constant D, nuclear spin-orbit interaction constant a, Fermi contact interaction constant b_{F}, magnetic dipolar interaction constant c, and nuclear quadrupole interaction constant eQq, were determined for the vibrationally ground state by least squares fitting of the observed spectrum. The equilibrium rotational constant B_{e} was determined to be 4435.7510(18) MHz and the internuclear distance between Co and C was evaluated to be 1.688 Å. The a and c values are consistent with the values estimated from the hyperfine constants of the ^{59}Co atom. The b_{F} value is nearly equal to zero.